OpenStudy (anonymous):

evaluate cot (cos^-1(-15/17))

OpenStudy (john_es):

Use the relation: \[1+\cot^2x=\frac{1}{1-\cos^2x}\]In this case, \[\cos x=-15/17\]So, \[1+\cot^2x=\frac{1}{1-(15/17)^2}\]Solve for cot x.